# Vendi-RAG: Adaptively Trading-Off Diversity And Quality Significantly Improves Retrieval Augmented Generation With LLMs

### v1

@@ Abstract
Retrieval-augmented generation (RAG) enhances large language models (LLMs) for domain-specific question-answering (QA) tasks by leveraging external knowledge sources. However, traditional RAG systems primarily focus on relevance-based retrieval and often struggle with redundancy, especially when reasoning requires connecting information from multiple sources. This paper introduces Vendi-RAG, a framework based on an iterative process that jointly optimizes retrieval diversity and answer quality. This joint optimization leads to significantly higher accuracy for multi-hop QA tasks. Vendi-RAG leverages the Vendi Score (VS), a flexible similarity-based diversity metric, to promote semantic diversity in document retrieval. It then uses an LLM judge that evaluates candidate answers, generated after a reasoning step, and outputs a score that the retriever uses to balance relevance and diversity among the retrieved documents during each iteration. Experiments on three challenging datasets—HotpotQA, MuSiQue, and 2WikiMultiHopQA—demonstrate Vendi-RAG’s effectiveness in multi-hop reasoning tasks. The framework achieves significant accuracy improvements over traditional single-step and multi-step RAG approaches, with accuracy increases reaching up to +4.2% on HotpotQA, +4.1% on 2WikiMultiHopQA, and +1.3% on MuSiQue compared to Adaptive-RAG, the current best baseline. The benefits of Vendi-RAG are even more pronounced as the number of retrieved documents increases. Finally, we evaluated Vendi-RAG across different LLM backbones, including GPT-3.5, GPT-4, and GPT-4o-mini, and observed consistent improvements, demonstrating that the framework’s advantages are model-agnostic.
@@ 1 Introduction
Retrieval-augmented generation (RAG) has emerged as a transformative framework for enhancing the performance of large language models (LLMs) in domain-specific tasks such as question-answering (QA). By retrieving relevant information from external sources beyond the training set, RAG enables LLMs to answer specialized queries more effectively (Achiam et al., 2023; Team et al., 2023; Jiang et al., 2024). This approach has been particularly successful in single-hop QA, where a question can be answered using information from a single document Raiaan et al. (2024); Kwiatkowski et al. (2019). For instance, answering a question such as "Which country is filmmaker Sembene Ousmane from?" only requires retrieving relevant information from a single document containing this fact.
@@
However, multi-hop QA introduces significantly greater complexity. Finding the correct answer to queries in multi-hop QA requires reasoning across multiple sources (Press et al., 2022; Tang and Yang, 2024). For instance, answering "Which city is the capital of the African country where Mount Kilimanjaro is located?" necessitates first identifying that Mount Kilimanjaro is in Tanzania, and then determining that Dodoma is the capital of Tanzania. This process involves not only retrieving information from multiple documents but also synthesizing these different sources effectively to form an accurate answer, which greatly increases the complexity of both retrieval and reasoning and often leads to redundancy.
@@
To address these challenges, iterative RAG pipelines have been developed. These pipelines refine the retrieval process through repeated modifications and re-querying of retrieved documents, aiming to resolve ambiguities and improve relevance. Notable examples include Adaptive-RAG (Jeong et al., 2024), Self-RAG (Asai et al., 2023), and IROC (Trivedi et al., 2022).
@@
Despite their success, iterative RAG methods typically rely solely on relevance-based retrieval, focusing on the similarity between the query and dataset entries and overlooking diversity. However, effectively addressing more complex queries requires diverse retrieval. We therefore propose a novel retrieval method called Vendi retrieval to address the limitation of existing retrieval pipelines. Vendi retrieval leverages the Vendi Score (VS) to enhance the diversity of retrieved documents while accounting for retrieval quality through a simple weighting mechanism.
@@
Building on Vendi retrieval, we propose an iterative RAG pipeline called Vendi-RAG that effectively balances diversity and quality. More specifically, an initial set of candidate documents is retrieved. Based on these retrieved documents, the system generates chain-of-thought (CoT) reasoning steps. Using these reasoning steps and retrieved documents, a backbone LLM then generates candidate answers. Another LLM, which we’ll refer to as an LLLM judge, then assesses these candidates for relevance, coherence, and completeness. The highest-scoring answer is selected as the final response. If the answer does not meet the quality threshold, the Vendi retrieval process dynamically adjusts the balance between diversity and relevance in document selection for the next iteration. This iterative refinement continues until a high-quality response is achieved. Figure 1 provides a detailed overview of the Vendi-RAG framework.
@@
We evaluated the Vendi retrieval process and Vendi-RAG on three challenging multi-hop QA datasets, HotpotQA (Yang et al., 2018), MuSiQue (Trivedi et al., 2022), and 2WikiMultiHopQA (Ho et al., 2020). To assess the Vendi retrieval method we measured the diversity of retrieved documents on the three datasets using two different diversity metrics, the VS and the max pairwise distance (MPD). We found that the Vendi retrieval process yields more diverse documents compared to the baselines according to both metrics. Second, we evaluated Vendi-RAG in terms of several performance metrics, looking at both accuracy and diversity. The results showed that Vendi-RAG substantially improves response accuracy, outperforming the baselines. Using GPT-3.5 as the LLM backbone, Vendi-RAG demonstrated significant accuracy gains across all datasets, with accuracy increases reaching +4.2% on HotpotQA, +4.1% on 2WikiMultiHopQA, and +1.3% on MuSiQue compared to Adaptive-RAG, the best baseline. Notably, the accuracy improvement remained consistent across different LLM backbones—GPT-4o, GPT-4o-mini, and GPT-3.5—indicating that Vendi-RAG’s advantages are model-agnostic. Additionally, our experiments with varying numbers of retrieved documents—beyond the standard two-document setting—showed that Vendi-RAG maintained its superior performance, especially as the number of retrieved documents increased. This underscores the critical role of the Vendi retrieval process in handling complex retrieval scenarios. For instance, when retrieving ten documents from HotpotQA, Vendi-RAG outperformed Adaptive-RAG by 7.8% in accuracy using GPT-4o-mini as the backbone LLM.
@@ 2 Related Work
Question answering.  There are three main approaches to QA: non-retrieval-based methods (Petroni et al., 2019), single-step RAG (Lewis et al., 2020), and multi-step RAG (Asai et al., 2023). Non-retrieval-based QA methods pass queries directly to an LLM and use its generated output as the answer, without consulting external sources. While efficient, these methods struggle with queries requiring external or up-to-date information and suffer from hallucinations on out-of-distribution queries (Shuster et al., 2021). Single-step RAG methods integrate external knowledge retrieved from a knowledge base (e.g., Wikipedia). These methods improve factual accuracy but are limited by retrieval noise and perform poorly in complex reasoning tasks (Trivedi et al., 2022). Multi-step RAG methods are designed for complex multi-hop queries (Jeong et al., 2024; Asai et al., 2023; Tang and Yang, 2024).
Recent improvements in multi-hop QA focus on question decomposition (Radhakrishnan et al., 2023), chain-of-thought reasoning (Wei et al., 2022; Liu et al., 2024a), and iterative retrieval (Jeong et al., 2024; Shao et al., 2023; Yu et al., 2024). Methods like ReCite (Sun et al., 2022) and IRCoT (Trivedi et al., 2022) refine retrieval with progressive reasoning, while Self-RAG (Asai et al., 2023) adapts retrieval strategies based on query complexity. Decomposed prompting (Khot et al., 2022) further enhances retrieval for complex queries (Zhang et al., 2024). MultiHop-RAG (Tang and Yang, 2024) integrates decomposition and retrieval pipelines but remains constrained by the redundancy issue caused by relevance-based retrieval.
Vendi scoring.  The Vendi Score (VS) (Friedman and Dieng, 2023) is a similarity-based diversity metric applied in machine learning (Berns et al., 2023; Pasarkar and Dieng, 2023; Mousavi and Khalili, 2025; Nguyen and Dieng, 2024; Kannen et al., 2024; Jalali et al., 2024; Askari Hemmat et al., 2024; Rezaei and Dieng, 2025; Bhardwaj et al., 2025), chemistry (Pasarkar et al., 2023), materials science (Liu et al., 2024b), and biology (Pasarkar and Dieng, 2025). Vendi-RAG integrates VS into retrieval, balancing diversity and quality beyond conventional ranking systems (Carbonell and Goldstein, 1998; Slivkins et al., 2010). Unlike standard relevance-based retrieval (Guu et al., 2020), this approach enhances robustness and accuracy in multi-hop QA by incorporating semantic diversity into document retrieval.
@@ 3.1 Vendi Retrieval
Diversity in retrieved documents is essential for multi-hop QA, as it ensures broad semantic coverage, reduces redundancy, and incorporates multiple perspectives (Sun et al., 2022; Carbonell and Goldstein, 1998; Thakur et al., 2021). The most used methods for diverse retrieval are similarity search (SS) (Thakur et al., 2021) and maximal marginal relevance (MMR) (Carbonell and Goldstein, 1998). SS maximizes relevance to the query but retrieves highly similar documents, leading to redundancy. MMR balances relevance and novelty using pairwise comparisons but also struggles to account for global semantic diversity.
To overcome these limitations, we propose a novel retrieval method that leverages the VS (Friedman and Dieng, 2023) to explicitly optimize retrieval diversity. Let $\mathcal{D}=\left\{d_{1},\dots,d_{n}\right\}$ be a set of retrieved documents and $k(\cdot,\cdot)$ a positive semi-definite similarity kernel such $k(d_{i},d_{i})=1$ for all $i$. Denote by K the corresponding similarity matrix that is such that $K_{ij}=k(d_{i},d_{j})$. The VS is defined as
@@
where $\lambda_{1},\dots,\lambda_{n}$ are the eigenvalues of the normalized kernel matrix $K/n$. As argued by Friedman and Dieng (2023), the VS is the effective number of unique documents in $\mathcal{D}$, reaching its maximum value $n$ when all the documents are distinct and its minimal value $1$ when all the documents are the same.
While accounting for diversity is good for retrieval, especially for complex queries, it shouldn’t be the only criterion. Quality also matters. To balance these two criteria, the Vendi retrieval process uses a convex combination of the two,
$\text{VRS}=s\cdot\text{VS}_{k}(q,\mathcal{D})+(1-s)\cdot\text{SS}(q,\mathcal{D}),$ (2)
where VRS stands for Vendi retrieval score. The similarity score $\text{SS}(q,\mathcal{D})$ is computed using dense vector representations of both the query and the documents. The document representations are used to provide meaningful context, ensuring that the retrieved documents are relevant to the query. Here $s\in[0,1]$ is a tunable parameter controlling the trade-off between diversity and similarity. When handling complex queries, such as those with multiple possible answers, a higher diversity weight $s$ promotes the selection of a semantically diverse set of documents. In contrast, for simpler or more specific queries that require precise information, a smaller value of $s$ prioritizes similarity-based retrieval.
@@ 3.2 Vendi-RAG
1. Initial retrieval.  The process begins by retrieving a set of documents using Vendi retrieval. This first step prioritizes broad semantic coverage (we set $s=0.8$ initially in all our experiments), ensuring that the retrieved documents capture multiple perspectives and to prevent recovering semantically redundant documents. This initial diversity is particularly critical for multi-hop QA, where synthesizing information from varied sources is essential to accurately answering the query.
2. Reasoning generation.  Based on the retrieved documents, the system generates CoT reasoning steps. These intermediate reasoning steps help contextualize the retrieved information, building a coherent pathway to the final answer.
3. Candidate answer generation.  Using the reasoning steps and retrieved documents, the LLM generates candidate answers. These proposed answers are evaluated to determine their quality and completeness.
4. Quality evaluation.  An LLM judge assesses the candidate answers. This evaluation considers factors such as coherence, relevance, and alignment with the query. A quality score $Q_{t}$ is produced at the end of this quality-check. Here $t$ is used to indicate the iteration step.
5. Dynamic adjustment of the VRS.  Based on the quality score $Q_{t}$, the parameter $s$ is adjusted dynamically. We denote by $s_{t}$ the value of the parameter $s$ at the $t^{\text{th}}$ iteration. It controls the trade-off between diversity (via VS) and relevance (via similarity search). If $Q_{t}$ is low, $s_{t}$ should be increased, to prioritize greater diversity in the subsequent retrievals. This ensures broader semantic exploration, which is beneficial for refining answers in cases where the retrieved information is already relevant but lacks coverage. Conversely, if $Q_{t}$ is high, $s_{t}$ should be decreased to focus more on relevance, retrieving documents that are closely aligned with the query to address potential gaps in specificity. We therefore define $s_{t}$ as
@@
6. Iterative refinement.  The retrieval and reasoning steps are repeated iteratively, with adjustments to $s$ dynamically balancing diversity and relevance at each stage. This process continues until the desired answer quality is reached, ensuring that the system converges on an optimal set of documents and reasoning steps.
7. Final answer selection.  Once the iterative refinement process is complete, the final set of documents and answers are selected based on their quality scores. This ensures that the output reflects both broad semantic coverage and high-quality, relevant information. Algorithm 1 summarizes the procedure.
Why Adjusting $s$ Matters:  The dynamic adjustment of $s$ is critical for striking the right balance between diversity and relevance during the retrieval process. High diversity is essential for exploring various facets of a complex query, especially in multi-hop QA, where information from disparate sources must be synthesized. However, excessive diversity can dilute the relevance of retrieved documents, potentially introducing noise and reducing the quality of generated answers. On the other hand, overemphasizing relevance can lead to redundancy and failure to capture the breadth of information needed for comprehensive reasoning.
By reducing $s$ when the quality score is high, the Vendi-RAG pipeline encourages exploration of less-redundant, semantically diverse documents. This ensures that even if the current answer is sufficient, the model explores additional perspectives that may enhance the depth and breadth of the final response. Conversely, increasing $s$ when quality is low allows the system to focus on retrieving documents that are more closely aligned with the query, addressing gaps in specificity or relevance.
The strength of Vendi-RAG lies in this adaptive approach to document retrieval. Unlike traditional RAG systems that use fixed retrieval strategies, Vendi-RAG’s dynamic adjustment of the diversity-relevance trade-off (the parameter $s$) allows it to respond to the specific requirements of each query and intermediate reasoning step. When the system detects that current retrievals are yielding high-quality but potentially narrow responses, it automatically shifts toward greater diversity, exploring complementary perspectives. Conversely, when responses lack precision, the system can focus on more closely related documents to improve specificity.
Performance characteristics.  In practice, Vendi-RAG exhibits distinctive performance patterns that reflect its sophisticated design. The system naturally adapts its computational effort to query complexity, requiring more iterations for intricate multi-hop queries while converging quickly for simpler ones. Though the computational overhead exceeds that of basic RAG systems, the improved retrieval quality often results in better final answers. The system maintains reasonable scalability with document corpus size, as the primary computational bottleneck—eigenvalue computation—depends on the number of retrieved documents rather than the total corpus size. These characteristics make Vendi-RAG particularly suitable for complex tasks such as multi-hop QA.
@@
      Compute scores: $\text{scores}_{i}=s_{i}\cdot\text{VS}_{k}(q,\mathcal{D})+(1-s_{i})\cdot\text{SS}(q,\mathcal{D})$
      Retrieve documents: $D_{i}\leftarrow\text{Vendi-Retriever}(q,\text{scores}_{i};\mathcal{D})$
@@
      Evaluate quality: $Q_{i}\leftarrow\text{Grader}(\hat{a}_{i})$
@@
return $\hat{a}^{*}$ // Return the best answer after $N$ iterations.

### v2

@@ Abstract
Retrieval-augmented generation (RAG) enhances large language models (LLMs) for domain-specific question-answering (QA) tasks by leveraging external knowledge sources. However, traditional RAG systems primarily focus on relevance-based retrieval and often struggle with redundancy, especially when reasoning requires connecting information from multiple sources. This paper introduces Vendi-RAG, a framework based on an iterative process that jointly optimizes retrieval diversity and answer quality. This joint optimization leads to significantly higher accuracy for multi-hop QA tasks. Vendi-RAG leverages the Vendi Score (VS), a flexible similarity-based diversity metric, to promote semantic diversity in document retrieval. It then uses an LLM judge that evaluates candidate answers, generated after a reasoning step, and outputs a score that the retriever uses to balance relevance and diversity among the retrieved documents during each iteration. Experiments on three challenging datasets—HotpotQA, MuSiQue, and 2WikiMultiHopQA—demonstrate Vendi-RAG’s effectiveness in multi-hop reasoning tasks. The framework achieves significant accuracy improvements over traditional single-step or multi-step RAG approaches, with accuracy increases reaching +4.2% on HotpotQA, +4.1% on 2WikiMultiHopQA, and +1.3% on MuSiQue compared to Adaptive-RAG, the current best baseline. The benefits of Vendi-RAG are even more pronounced as the number of retrieved documents increases. Finally, we evaluated Vendi-RAG across different LLM backbones, including GPT-3.5, GPT-4, and GPT-4o-mini, and observed consistent improvements, demonstrating that the framework’s advantages are model-agnostic.
@@ 1 Introduction
Retrieval-augmented generation (RAG) has emerged as a transformative framework for enhancing the performance of large language models (LLMs) in domain-specific tasks such as question-answering (QA). By retrieving relevant information from external sources beyond the training set, RAG enables LLMs to answer specialized queries more effectively Achiam et al. (2023); Team et al. (2023); Jiang et al. (2024). This approach has been particularly successful in single-hop QA, where a question can be answered using information from a single document Raiaan et al. (2024); Kwiatkowski et al. (2019). For instance, answering a question such as "Who wrote the novel Frankenstein?" only requires retrieving relevant information from a single document containing this fact.
@@
However, multi-hop QA introduces significantly greater complexity. Finding the correct answer to queries in multi-hop QA requires reasoning across multiple sources (Press et al., 2022; Tang and Yang, 2024). For instance, answering "Which city is the capital of the African country where Mount Kilimanjaro is located?" necessitates first identifying that Mount Kilimanjaro is in Tanzania, and then determining that Dodoma is the capital of Tanzania. This process involves not only retrieving information from multiple documents but also synthesizing these different sources effectively to form an accurate answer, which greatly increases the complexity of both retrieval and reasoning and leads to redundancy.
@@
To address these challenges, iterative RAG pipelines have been developed. These pipelines refine the retrieval process through repeated modifications and re-querying of retrieved documents, aiming to resolve ambiguities and improve relevance. Notable examples include Adaptive-RAG (Lewis et al., 2020), which controls the number of iterations of the pipeline including the retrieval process and modifying the queries based on a classification model’s assessment of the input query, Self-RAGAsai et al. (2023), which incorporates iterative self-reasoning, and IROC Trivedi et al. (2022), which progressively refines retrieval to optimize the final answer Wei et al. (2022); Wang and Zhou (2024).
@@
Despite their success, iterative RAG methods typically rely solely on relevance-based retrieval, which focuses on the similarity between the query and dataset entries. This approach presents a fundamental limitation: it does not actively manage the diversity and quality of the retrieved information to properly address the query. More complex queries require diverse retrieval. We therefore propose a novel retrieval method called Vendi retrieval to address the limitation of existing retrieval pipelines. Vendi retrieval leverages the Vendi Score (VS) to enhance the diversity of retrieved documents while accounting for retrieval quality through a simple weighting mechanism.
@@
Building on Vendi retrieval, we propose an iterative RAG pipeline called Vendi-RAG that balances diversity and quality. More specifically, the pipeline is as follows: an initial set of candidate documents is retrieved. Based on these retrieved documents, the system generates chain-of-thought (CoT) reasoning steps. Using these reasoning steps and retrieved documents, the LLM then generates candidate answers. An LLM-based evaluator then assesses these candidates for relevance, coherence, and completeness. The highest-scoring answer is selected as the final response. If the answer does not meet the quality threshold, the Vendi retrieval process dynamically adjusts the balance between diversity and relevance in document selection, ensuring broader semantic exploration or increased specificity as needed. This iterative refinement continues until a high-quality response is achieved. Figure 1 provides a detailed overview of the Vendi-RAG framework.
@@
We evaluated the Vendi retrieval process and Vendi-RAG on three challenging multi-hop QA datasets, HotpotQA (Yang et al., 2018), MuSiQue (Trivedi et al., 2022), and 2WikiMultiHopQA (Ho et al., 2020). To assess the Vendi retrieval method we measured the diversity of retrieved documents on the three datasets using two different diversity metrics, the VS and the max pairwise distance (MPD). We found that the Vendi retrieval process yields more diverse documents compared to the baselines according to both metrics. Second, we evaluated Vendi-RAG in terms of several performance metrics, looking at both accuracy and diversity. The results showed that Vendi-RAG substantially improves response accuracy, outperforming existing RAG approaches. Using GPT-3.5 as the LLM backbone, Vendi-RAG demonstrated significant accuracy gains across all datasets, with accuracy increases reaching +4.2% on HotpotQA, +4.1% on 2WikiMultiHopQA, and +1.3% on MuSiQue compared to Adaptive-RAG, the best baseline. Notably, the accuracy improvement remained consistent across different LLM backbones—GPT-4o, GPT-4o-mini, and GPT-3.5—indicating that Vendi-RAG’s advantages are model-agnostic. Additionally, our experiments with varying numbers of retrieved documents—beyond the standard two-document setting—showed that Vendi-RAG maintained its superior performance, especially as the number of retrieved documents increased. This underscores the critical role of the Vendi retrieval process in handling complex retrieval scenarios. For instance, when retrieving ten documents from HotpotQA, Vendi-RAG outperformed Adaptive-RAG by 7.8% in accuracy using GPT-4o-mini as the backbone LLM.
This work introduces a diversity-guided retrieval approach that optimizes both diversity and quality to address the challenges of multi-step reasoning in multi-hop QA. Our experimental results highlight the effectiveness of Vendi-RAG in enhancing retrieval diversity and response accuracy, underscoring its potential as a robust solution for complex multi-hop QA tasks.
@@ 2 Related Work
There are three main approaches to QA: non-retrieval-based methods (Petroni et al., 2019), single-step RAG (Lewis et al., 2020), and multi-step RAG (Asai et al., 2023). Non-retrieval-based QA methods pass queries directly to an LLM and use its generated output as the answer, without consulting external sources. While efficient, these methods struggle with queries requiring external or up-to-date information and suffer from hallucinations on out-of-distribution queries (Shuster et al., 2021). Single-step RAG methods integrate external knowledge retrieved from a knowledge base (e.g., Wikipedia). These methods improve factual accuracy but are limited by retrieval noise and perform poorly in complex reasoning tasks (Trivedi et al., 2022). Multi-step RAG methods are designed for complex multi-hop queries (Jeong et al., 2024; Asai et al., 2023; Tang and Yang, 2024). They iteratively retrieve documents and refine answers until they converge on a final response. This iterative refinement approach enables reasoning across multiple sources but introduces computational overhead and is prone to error accumulation (Jeong et al., 2024).
Advances in multi-hop QA.
Recent improvements in multi-hop QA focus on question decomposition (Radhakrishnan et al., 2023), chain-of-thought reasoning (Wei et al., 2022; Liu et al., 2024a), and iterative retrieval (Jeong et al., 2024; Shao et al., 2023; Yu et al., 2024). Methods like ReCite (Sun et al., 2022) and IRCoT (Trivedi et al., 2022) refine retrieval with progressive reasoning, while Self-RAG (Asai et al., 2023) adapts retrieval strategies based on query complexity. Decomposed prompting (Khot et al., 2022) further enhances retrieval for complex queries (Zhang et al., 2024). MultiHop-RAG (Tang and Yang, 2024) integrates decomposition and retrieval pipelines but remains constrained by relevance-based retrieval, leading to redundancy and limited synthesis of diverse information.
Vendi scoring.
The Vendi Score (VS) (Friedman and Dieng, 2023) is a similarity-based diversity metric applied in machine learning (Berns et al., 2023; Pasarkar and Dieng, 2023; Mousavi and Khalili, 2025; Nguyen and Dieng, 2024; Kannen et al., 2024; Jalali et al., 2024; Askari Hemmat et al., 2024; Rezaei and Dieng, 2025; Bhardwaj et al., 2025), chemistry (Pasarkar et al., 2023), materials science (Liu et al., 2024b), and biology (Pasarkar and Dieng, 2025). Vendi-RAG integrates VS into retrieval, balancing diversity and quality beyond conventional ranking systems (Carbonell and Goldstein, 1998; Slivkins et al., 2010). Unlike standard relevance-based retrieval (Guu et al., 2020), this approach enhances robustness and accuracy in multi-hop QA by incorporating semantic diversity into document retrieval.
@@ 3.1 Vendi Retrieval
Diversity in retrieved documents is essential for multi-hop QA, as it ensures broad semantic coverage, reduces redundancy, and incorporates multiple perspectives (Sun et al., 2022; Carbonell and Goldstein, 1998; Thakur et al., 2021). The most commonly used techniques for information retrieval are similarity search (SS)(Thakur et al., 2021) and maximal marginal relevance (MMR)(Carbonell and Goldstein, 1998). While SS retrieves documents based on their relevance to the query, it often results in redundant documents with high similarity. MMR attempts to balance relevance and novelty using pairwise comparisons, but it still struggles to capture global semantic diversity.
To address these limitations, we adopt a retrieval approach based on the Vendi Score (VS) (Friedman and Dieng, 2023), which explicitly quantifies semantic diversity in a set of documents. Let $\mathcal{D}={d_{1},\dots,d_{n}}$ be a set of retrieved documents, and let $k(\cdot,\cdot)$ be a positive semi-definite similarity kernel such that $k(d_{i},d_{i})=1$ for all $i$. Let $K$ be the similarity matrix with entries $K_{ij}=k(d_{i},d_{j})$. The Vendi Score is
@@
where $\lambda_{1},\dots,\lambda_{n}$ are the eigenvalues of the normalized kernel matrix $K/n$. As shown by Friedman and Dieng (2023), $\text{VS}_{k}(\mathcal{D})$ reflects the effective number of unique documents in $\mathcal{D}$, attaining its maximum value $n$ when all documents are orthogonal (fully diverse) and its minimum value 1 when all documents are identical.
While optimizing for diversity is important—especially for complex, multi-faceted queries—it must be balanced with query relevance. To achieve this, we define the Vendi Retrieval Score (VRS) as a convex combination of semantic diversity and similarity-based relevance:
$\text{VRS}=s\cdot\text{VS}_{k}(\mathcal{D})+(1-s)\cdot\text{SS}(q,\mathcal{D}),$ (2)
where $s\in[0,1]$ is a tunable parameter that controls the trade-off between diversity and relevance. The similarity score $\text{SS}(q,\mathcal{D})$ is computed using dense vector representations of the query $q$ and the documents in $\mathcal{D}$, typically obtained from transformer-based encoders. This ensures semantic matching beyond surface-level lexical overlap.
It is important to note that while the Vendi Score $\text{VS}_{k}(\mathcal{D})$ is computed solely based on the retrieved documents and their pairwise similarities, query relevance is introduced in the initial retrieval step: the candidate set $\mathcal{D}$ is selected using similarity search with respect to the query $q$. Thus, the formulation in Equation (2) balances document-level diversity and query-level relevance, where a higher value of $s$ favors diverse content, and a lower value prioritizes semantic similarity to the query. In this way, the VSR addresses the dual objectives of reducing redundancy and maintaining relevance, providing a principled and flexible framework for multi-hop document selection.
@@ 3.2 Vendi-RAG
1. Initial retrieval.
The process begins by retrieving a set of documents using Vendi retrieval. This first step prioritizes broad semantic coverage (we set $s=0.8$ initially in all our experiments), ensuring that the retrieved documents capture multiple perspectives and to prevent recovering semantically redundant documents. This initial diversity is particularly critical for multi-hop QA, where synthesizing information from varied sources is essential to accurately answering the query.
2. Reasoning generation.
Based on the retrieved documents, the system generates CoT reasoning steps. These intermediate reasoning steps help contextualize the retrieved information, building a coherent pathway to the final answer.
3. Candidate answer generation.
Using the reasoning steps and retrieved documents, the LLM generates candidate answers. These proposed answers are evaluated to determine their quality and completeness.
4. Quality evaluation.
An LLM judge assesses the candidate answers. This evaluation considers factors such as coherence, relevance, and alignment with the query. A quality score $Q_{t}$ is produced at the end of this quality-check. Here $t$ is used to indicate the iteration step.
5. Dynamic adjustment of the VRS.
Based on the quality score $Q_{t}$, the parameter $s$ is adjusted dynamically. We denote by $s_{t}$ the value of the parameter $s$ at the $t^{\text{th}}$ iteration. It controls the trade-off between diversity (via VS) and relevance (via similarity search). If $Q_{t}$ is low, $s_{t}$ should be increased, to prioritize greater diversity in the subsequent retrievals. This ensures broader semantic exploration, which is beneficial for refining answers in cases where the retrieved information is already relevant but lacks coverage. Conversely, if $Q_{t}$ is high, $s_{t}$ should be decreased to focus more on relevance, retrieving documents that are closely aligned with the query to address potential gaps in specificity. We therefore define $s_{t}$ as
@@
6. Iterative refinement.
The retrieval and reasoning steps are repeated iteratively, with adjustments to $s$ dynamically balancing diversity and relevance at each stage. This process continues until the desired answer quality is reached, ensuring that the system converges on an optimal set of documents and reasoning steps.
7. Final answer selection.
Once the iterative refinement process is complete, the final set of documents and answers are selected based on their quality scores. This ensures that the output reflects both broad semantic coverage and high-quality, relevant information. Algorithm 1 summarizes the procedure.
Why Adjusting $s$ Matters:
The dynamic adjustment of $s$ is essential for balancing diversity and relevance during retrieval. High diversity enables exploration of different facets of complex queries, especially in multi-hop QA, where synthesizing information from multiple sources is crucial. However, too much diversity can introduce noise, while excessive focus on relevance risks redundancy and limits comprehensive reasoning.
Vendi-RAG addresses this by adapting $s$ based on retrieval quality: when the quality score is high, it reduces $s$ to promote exploration of additional, semantically diverse documents; when quality is low, it increases $s$ to prioritize more directly relevant documents. This adaptive retrieval strategy allows Vendi-RAG to dynamically adjust to the needs of each query and reasoning step, improving both the breadth and precision of generated answers. Unlike traditional RAG systems with fixed retrieval policies, Vendi-RAG’s flexibility ensures richer, more contextually appropriate responses.
Performance characteristics.
In practice, Vendi-RAG exhibits distinctive performance patterns that reflect its sophisticated design. The system naturally adapts its computational effort to query complexity, requiring more iterations for intricate multi-hop queries while converging quickly for simpler ones. Though the computational overhead exceeds that of basic RAG systems, the improved retrieval quality often results in better final answers. The system maintains reasonable scalability with document corpus size, as the primary computational bottleneck—eigenvalue computation—depends on the number of retrieved documents rather than the total corpus size. These characteristics make Vendi-RAG particularly suitable for complex tasks such as multi-hop QA.
@@
      Compute scores: $\text{VSR}_{i}=s_{i}\cdot\text{VS}_{k}(\mathcal{D})+(1-s_{i})\cdot\text{SS}(q,\mathcal{D})$
      Get the document set: $D_{i}\leftarrow\text{Vendi-Retriever}(\text{VSR}_{i};\mathcal{D})$
@@
      Evaluate quality: $Q_{i}\leftarrow\text{LLM-Judge}(\hat{a}_{i})$
@@
return $\hat{a}_{N}$ // Return the best answer after $N$ iterations.